What is the percentage concentration of 75mL of ethanol dissolved in 500mL of water?





1
Expert's answer
2021-11-08T08:11:53-0500

Solution:

solute = ethanol (75 mL)

solvent = water (500 mL)

volume of solution = volume of solute + volume of solvent = 75 mL + 500 mL = 575 mL


Percent concentration by volume, (v/v)%:

(v/v)% = (volume of solute / volume of solution) × 100%

Hence,

(v/v)% = (75 mL / 575 mL) × 100% = 13.043% = 13%

(v/v)% = 13% ethanol


Answer: The percentage concentration by volume of ethanol is 13%.